# A Gaussian Process View on Observation Noise and Initialization in Wide Neural Networks

### v2

@@ D.1 Regularized gradient flow for the linearized network
$f^{\operatorname*{lin}}(\mathbf{x^{\prime}},\theta^{\operatorname*{lin}}_{t})$ $=f(\mathbf{x^{\prime}},\theta_{0})+J(\mathbf{x^{\prime}},\theta_{0})J(\theta_{0})^{\top}\left(I_{N}-e^{-\eta_{0}\left(J(\theta_{0})J(\theta_{0})^{\top}+\beta I_{N}\right)t}\right)\left(J(\theta_{0})J(\theta_{0})^{\top}+\beta I_{N}\right)^{-1}(\mathbf{y}-f(\theta_{0}))$ (41)
@@ D.2 Regularized gradient descent for the linearized network
$f^{\operatorname*{lin}}(\mathbf{x^{\prime}},\theta^{\operatorname*{lin}}_{t})$ $=f(\mathbf{x^{\prime}},\theta_{0})+J(\mathbf{x^{\prime}},\theta_{0})J(\theta_{0})^{\top}\left(I_{N}-\left(I_{N}-\eta_{0}\left(J(\theta_{0})J(\theta_{0})^{\top}+\beta I_{N}\right)\right)^{t}\right)\left(J(\theta_{0})J(\theta_{0})^{\top}+\beta I_{N}\right)^{-1}(\mathbf{y}-f(\theta_{0}))$ (49)
@@ E.1 Convergence of NTK under NTK parametrization at initialization
Here, we restate the following Theorem from yang2020tensorprograms2 about the convergence of the NTK at initialization. This was first shown in jacot2018neural when taking the limit of layer widths sequentially.
@@
Define $\mathbf{\Theta}:=\mathbf{\Theta}_{\mathbf{x},\mathbf{x}}\in\mathbb{R}^{N\times N}$ as the analytical NTK on the training points. We will assume $\lambda_{\min}(\mathbf{\Theta})>0$. A sufficient condition for this is that $\lVert\mathbf{x}_{i}\rVert_{2}=1\;\forall i$, and that $\phi$ grows non-polynomially for large $x$, see jacot2018neural. This directly implies that for $n$ large enough, the minimum eigenvalue of the analytical NTK is lower bounded by a positive number with high probability:
@@ E.3 Redundancies when using the same learning rate for standard parametrization
The learning rate that was used in the proof of lee2019wide is $\frac{1}{n}\eta_{0}$ for any layer. In the following, we will argue that this effectively leads to the first layer, and the biases not being trained in the infinite-width limit. For simplicity, let $\beta=0$. Remember that lee2019wide shows that using the learning rate $\frac{1}{n}\eta_{0}$ for each layer in standard parametrization, leads to the trained network for large width being driven by the standard parametrization NTK
@@
Thus, the analytical standard parametrization NTK of lee2019wide doesn’t depend on the contribution of the gradient with respect to the first layer and the biases. In other words, using the learning rate $\frac{1}{n}\eta_{0}$ for the first layer and the biases leads to them not being trained for large widths.
@@
$J^{\operatorname*{std}}(\theta)HJ^{\operatorname*{std}}(\theta)=J^{\operatorname*{std}}(W^{1})J^{\operatorname*{std}}(W^{1})^{\top}+J^{\operatorname*{std}}(b^{1})J^{\operatorname*{std}}(b^{1})^{\top}+\sum_{l=2}^{L+1}\left(\frac{1}{n}J^{\operatorname*{std}}(W^{l})J^{\operatorname*{std}}(W^{l})^{\top}+J^{\operatorname*{std}}(b^{l})J^{\operatorname*{std}}(b^{l})^{\top}\right).$ (70)
For simplicity, we will not train the first layer and the biases.
@@ F.1 Exponential Decay of the Regularized Gradient and Closeness of Parameters to their Initial Value
Hence, $\lVert g(\theta_{t})\rVert_{2}\leq\lVert g(\theta_{0})\rVert_{2}$. Further, note that $f(\theta_{0})$ converges in distribution to a Gaussian with mean zero and covariance given by the NNGP kernel (lee2018dnnsgps). Thus, for $n$ large enough, one can bound $\lVert g(\theta_{0})\rVert_{2}$ with high probability. ∎
@@
We want to show that the gradient $\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t})$ of the regularized loss converges to 0 quickly, and hence $\theta_{t}$ doesn’t move much. For $\beta=0$, its norm is $\lVert J(\theta_{t})^{\top}g(\theta_{t})\rVert_{2}$ and hence lee2019wide and related proofs showed that $\lVert g(\theta_{t})\rVert_{2}$ converges to 0 quickly. However, for $\beta>0$, this is not the case, as the training error won’t converge to 0. Instead, we directly look at the dynamics of the norm of the gradient:
@@
Using Lemma F.1, we have $\lVert g(\theta_{t})\rVert_{2}\leq\lVert g(\theta_{0})\rVert_{2}\leq R_{0}$. Further, by Lemma A.1, we have $\lVert\nabla_{\theta}^{2}f(\mathbf{x}_{i},\theta_{t})\rVert_{2}\leq\frac{1}{\sqrt{n}}K^{\prime}$. Thus (with $K=\sqrt{N}K^{\prime}$),
$\lVert g(\theta_{t})^{\top}\nabla_{\theta}^{2}f(\theta)\rVert_{2}\leq\frac{1}{\sqrt{n}}KR_{0}.$ (78)
@@
$\lambda_{\min}\left(g(\theta_{t})^{\top}\nabla_{\theta}^{2}f(\theta)\right)\geq-\lVert g(\theta_{t})^{\top}\nabla_{\theta}^{2}f(\theta)\rVert_{2}\geq-\frac{1}{\sqrt{n}}KR_{0}.$ (79)
@@
$\lambda_{\min}\left(\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t})\right)\geq-\frac{1}{\sqrt{n}}KR_{0}+0+\beta\geq\frac{\beta}{2}=:c_{\beta}.$ (80)
@@
Thus, for $t\leq t_{1}$, $\lVert\theta_{t}-\theta_{0}\rVert_{2}<C$. By continuity, $t_{1}=\infty$. Using local Lipschitzness, we can further bound the distance of the Jacobian at any $\lVert x\rVert_{2}\leq 1$
$\lVert J(x,\theta_{t})-J(x,\theta_{0})\rVert_{2}\leq\frac{1}{\sqrt{n}}K^{\prime}\lVert\theta_{t}-\theta_{0}\rVert_{2}\leq\frac{1}{\sqrt{n}}K^{\prime}C.$ (85)
This finishes the proof of Theorem 4.2.
@@
$\frac{d}{dt}\lVert\nabla_{\theta}\mathcal{L}^{0}(\theta_{t})\rVert_{2}^{2}$ $=-3\eta_{0}\left(\nabla_{\theta}\mathcal{L}^{0}(\theta_{t})\right)^{\top}\nabla_{\theta}^{2}\mathcal{L}^{0}(\theta_{t})\left(\nabla_{\theta}\mathcal{L}^{0}(\theta_{t})\right)$ (87)
$\leq-2\eta_{0}\lambda_{\min}\left(\nabla_{\theta}\mathcal{L}^{0}(\theta_{t})\right)\lVert\nabla_{\theta}\mathcal{L}^{0}(\theta_{t})\rVert_{2}^{2}-\eta_{0}\frac{1}{2}\lambda_{\min}(\mathbf{\Theta})\lVert\nabla_{\theta}\mathcal{L}^{0}(\theta_{t})\rVert_{2}^{2}$ (88)
$\leq-2\eta_{0}\left(-\frac{\log n}{\sqrt{n}}KR_{0}+\frac{1}{2}\lambda_{\min}(\mathbf{\Theta})\right)\lVert\nabla_{\theta}\mathcal{L}^{0}(\theta_{t})\rVert_{2}^{2}$ (89)
@@
Defining $c_{0}:=\frac{1}{3}\lambda_{\min}(\mathbf{\Theta})$ makes it possible to continue with the proof above for $\beta\geq 0$. This is an alternative proof to lee2019wide. It shows that in the unregularized case, it is important that the gradient flow lies in the row space of the Jacobian.
@@ F.2 Closeness to the Linearized Network along the Regularized Gradient Flow
The proof of lee2019wide used that training error converges to 0. Thus, we have to use a different approach, which also provides a more straightforward and intuitive proof for $\beta=0$. Remember that
@@
$\leq\frac{1}{\sqrt{n}}K^{\prime}C\eta_{0}KR_{0}e^{-\eta_{0}c_{\beta}t},$ (95)
@@
$\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}(x,\theta_{t})\rVert_{2}\leq\frac{1}{\sqrt{n}}K^{\prime}C\eta_{0}KR_{0}\int_{0}^{t}e^{-\eta_{0}c_{\beta}u}du\leq\frac{1}{\sqrt{n}}K^{\prime}C\frac{KR_{0}}{c_{\beta}}=\frac{1}{\sqrt{n}}K^{\prime}C^{2}.$ (96)
In particular, for the difference at the training points, $\lVert f(\theta_{t})-f^{\operatorname*{lin}}(\theta_{t})\rVert_{2}\leq\frac{1}{\sqrt{n}}KC^{2}$.
@@
$\lVert\left(J(\theta_{t})-J(\theta_{0})\right)^{\top}g(\theta_{t})\rVert_{2}\leq\lVert J(\theta_{t})-J(\theta_{0})\rVert_{2}\lVert g(\theta_{t})\rVert_{2}\leq\frac{1}{\sqrt{n}}KCR_{0}.$ (100)
@@
$\lVert J(\theta_{0})^{\top}\left(g(\theta_{t})-g^{\operatorname*{lin}}(\theta_{t})\right)\rVert_{2}=\lVert J(\theta_{0})^{\top}\left(f(\theta_{t})-f^{\operatorname*{lin}}(\theta_{t})\right)\rVert_{2}\leq\lVert J(\theta_{0})\rVert_{2}\lVert f(\theta_{t})-f^{\operatorname*{lin}}(\theta_{t})\rVert_{2}\leq\frac{1}{\sqrt{n}}K^{2}C^{2}.$ (101)
Thus, defining $K^{\Delta}:=KCR_{0}+K^{2}C^{2}$, we can bound $\lVert\Delta_{t}\rVert_{2}\leq\frac{1}{\sqrt{n}}K^{\Delta}$. Now, we can compute
$\frac{d}{dt}(\theta_{t}-\theta^{\operatorname*{lin}}_{t})$ $=-\eta_{0}\Delta_{t}-\eta_{0}\left(J(\theta_{0})^{\top}g^{\operatorname*{lin}}(\theta_{t})+\beta_{N}(\theta_{t}-\theta_{0})\right)+\eta_{0}\left(J(\theta_{0})^{\top}g^{\operatorname*{lin}}(\theta^{\operatorname*{lin}}_{t})+\beta_{N}(\theta^{\operatorname*{lin}}_{t}-\theta_{0})\right)$ (102)
@@
$\leq\int_{0}^{t}e^{-\eta_{0}(c_{0}+\beta)(t-u)}\eta_{0}\frac{1}{\sqrt{n}}K^{\Delta}du$ (108)
$\leq\frac{1}{\sqrt{n}}\frac{K^{\Delta}}{c_{0}+\beta}.$ (109)
Thus, $\sup_{t}\lVert\theta_{t}-\theta^{\operatorname*{lin}}_{t}\rVert_{2}\leq\frac{K^{\Delta}}{c_{0}+\beta}\frac{1}{\sqrt{n}}$.
@@
$\lVert f^{\operatorname*{lin}}(x,\theta_{t})-f^{\operatorname*{lin}}(x,\theta^{\operatorname*{lin}}_{t})\rVert_{2}=\lVert J(x,\theta_{0})(\theta_{t}-\theta^{\operatorname*{lin}}_{t})\rVert_{2}\leq\lVert J(x,\theta_{0})\rVert_{2}\lVert\theta_{t}-\theta^{\operatorname*{lin}}_{t}\rVert_{2}\leq K^{\prime}\frac{K^{\Delta}}{c_{0}+\beta}\frac{1}{\sqrt{n}}.$ (110)
By using Equation 92, we can now finish the proof:
$\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}(x,\theta^{\operatorname*{lin}}_{t})\rVert_{2}\leq\left(K^{\prime}C^{2}+K^{\prime}\frac{K^{\Delta}}{c_{0}+\beta}\right)\frac{1}{\sqrt{n}}.$ (111)

### v3

@@ D.1 Regularized gradient flow for the linearized network
$f^{\operatorname*{lin}}(\mathbf{x^{\prime}},\theta^{\operatorname*{lin}}_{t})$ (38)
$=$ $f(\mathbf{x^{\prime}},\theta_{0})+J(\mathbf{x^{\prime}},\theta_{0})J(\theta_{0})^{\top}\left(I_{N}-e^{-\eta_{0}\left(J(\theta_{0})J(\theta_{0})^{\top}+\beta I_{N}\right)t}\right)\left(J(\theta_{0})J(\theta_{0})^{\top}+\beta I_{N}\right)^{-1}(\mathbf{y}-f(\theta_{0}))$ (39)
@@ D.2 Regularized gradient descent for the linearized network
$f^{\operatorname*{lin}}(\mathbf{x^{\prime}},\theta^{\operatorname*{lin}}_{t})$ (47)
$=$ $f(\mathbf{x^{\prime}},\theta_{0})+J(\mathbf{x^{\prime}},\theta_{0})J(\theta_{0})^{\top}\left(I_{N}-\left(I_{N}-\eta_{0}\left(J(\theta_{0})J(\theta_{0})^{\top}+\beta I_{N}\right)\right)^{t}\right)\left(J(\theta_{0})J(\theta_{0})^{\top}+\beta I_{N}\right)^{-1}(\mathbf{y}-f(\theta_{0}))$ (48)
@@ E.1 Convergence of NTK under NTK parametrization at initialization
Here, we restate the following Theorem from Yang (2020) about the convergence of the NTK at initialization. This was first shown in Jacot et al. (2018) when taking the limit of layer widths sequentially.
@@
Define $\mathbf{\Theta}:=\mathbf{\Theta}_{\mathbf{x},\mathbf{x}}\in\mathbb{R}^{N\times N}$ as the analytical NTK on the training points. We will assume $\lambda_{\min}(\mathbf{\Theta})>0$. A sufficient condition for this is that $\lVert\mathbf{x}_{i}\rVert_{2}=1\;\forall i$, and that $\phi$ grows non-polynomially for large $x$, see Jacot et al. (2018). This directly implies that for $n$ large enough, the minimum eigenvalue of the analytical NTK is lower bounded by a positive number with high probability:
@@ E.3 Redundancies when using the same learning rate for standard parametrization
The learning rate that was used in the proof of Lee et al. (2019) is $\frac{1}{n}\eta_{0}$ for any layer. In the following, we will argue that this effectively leads to the first layer, and the biases not being trained in the infinite-width limit. For simplicity, let $\beta=0$. Remember that Lee et al. (2019) shows that using the learning rate $\frac{1}{n}\eta_{0}$ for each layer in standard parametrization, leads to the trained network for large width being driven by the standard parametrization NTK
@@
Thus, the analytical standard parametrization NTK of Lee et al. (2019) doesn’t depend on the contribution of the gradient with respect to the first layer and the biases. In other words, using the learning rate $\frac{1}{n}\eta_{0}$ for the first layer and the biases leads to them not being trained for large widths.
@@
$J^{\operatorname*{std}}(\theta)HJ^{\operatorname*{std}}(\theta)=$ $J^{\operatorname*{std}}(W^{1})J^{\operatorname*{std}}(W^{1})^{\top}+J^{\operatorname*{std}}(b^{1})J^{\operatorname*{std}}(b^{1})^{\top}$ (69)
$+\sum_{l=2}^{L+1}\left(\frac{1}{n}J^{\operatorname*{std}}(W^{l})J^{\operatorname*{std}}(W^{l})^{\top}+J^{\operatorname*{std}}(b^{l})J^{\operatorname*{std}}(b^{l})^{\top}\right).$ (70)
@@ F.1 Exponential Decay of the Regularized Gradient and Closeness of Parameters to their Initial Value
Hence, $\lVert g(\theta_{t})\rVert_{2}\leq\lVert g(\theta_{0})\rVert_{2}$. Further, note that $f(\theta_{0})$ converges in distribution to a Gaussian with mean zero and covariance given by the NNGP kernel (Lee et al., 2018). Thus, for $n$ large enough, one can bound $\lVert g(\theta_{0})\rVert_{2}$ with high probability. ∎
The proof of this Lemma does not apply to the weight-decay regularizer $\frac{1}{2}\lVert\theta\rVert_{2}^{2}$ analyzed in Lewkowycz and Gur-Ari (2020): In general, $\frac{1}{2}\lVert\theta_{0}\rVert_{2}^{2}\neq 0$, and thus we can not use $\mathcal{L}^{\beta}(\theta_{0})=\frac{1}{2}\lVert g(\theta_{0})\rVert_{2}^{2}$. They show that using gradient flow with this regularizer leads to $f_{\theta_{\infty}}(\cdot)=0$ in the infinite-width limit, and we are thus not in the NTK regime.
@@
We want to show that the gradient $\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t})$ of the regularized loss converges to 0 quickly, and hence $\theta_{t}$ doesn’t move much. For $\beta=0$, its norm is $\lVert J(\theta_{t})^{\top}g(\theta_{t})\rVert_{2}$ and hence Lee et al. (2019) and related proofs showed that $\lVert g(\theta_{t})\rVert_{2}$ converges to 0 quickly. However, for $\beta>0$, this is not the case, as the training error won’t converge to 0. Instead, we directly look at the dynamics of the norm of the gradient:
@@
Using Lemma F.1, we have $\lVert g(\theta_{t})\rVert_{2}\leq\lVert g(\theta_{0})\rVert_{2}\leq R_{0}$. Further, by Lemma A.1, we have $\lVert\nabla_{\theta}^{2}f(\mathbf{x}_{i},\theta_{t})\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}$. Thus (with $K=\sqrt{N}K^{\prime}$),
$\lVert g(\theta_{t})^{\top}\nabla_{\theta}^{2}f(\theta)\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}KR_{0}.$ (78)
@@
$\lambda_{\min}\left(g(\theta_{t})^{\top}\nabla_{\theta}^{2}f(\theta)\right)\geq-\lVert g(\theta_{t})^{\top}\nabla_{\theta}^{2}f(\theta)\rVert_{2}\geq-\frac{(\log n)^{c}}{\sqrt{n}}KR_{0}.$ (79)
@@
$\lambda_{\min}\left(\nabla_{\theta}\mathcal{L}^{\beta}(\theta_{t})\right)\geq-\frac{(\log n)^{c}}{\sqrt{n}}KR_{0}+0+\beta\geq\frac{\beta}{2}=:c_{\beta}.$ (80)
@@
Thus, for $t\leq t_{1}$, $\lVert\theta_{t}-\theta_{0}\rVert_{2}<C$. By continuity, $t_{1}=\infty$. Using local Lipschitzness, we can further bound the distance of the Jacobian at any $\lVert x\rVert_{2}\leq 1$:
$\lVert J(x,\theta_{t})-J(x,\theta_{0})\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}\lVert\theta_{t}-\theta_{0}\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}C.$ (85)
This finishes the proof of Theorem 4.2. ∎
@@
$\frac{d}{dt}\lVert\nabla_{\theta}\mathcal{L}^{0}(\theta_{t})\rVert_{2}^{2}$ $=-2\eta_{0}\left(\nabla_{\theta}\mathcal{L}^{0}(\theta_{t})\right)^{\top}\nabla_{\theta}^{2}\mathcal{L}^{0}(\theta_{t})\left(\nabla_{\theta}\mathcal{L}^{0}(\theta_{t})\right)$ (87)
$\leq-2\eta_{0}\left(\lambda_{\min}\left(g(\theta_{t})^{\top}\nabla_{\theta}^{2}f(\theta)\right)+\frac{1}{2}\lambda_{\min}(\mathbf{\Theta})\right)\lVert\nabla_{\theta}\mathcal{L}^{0}(\theta_{t})\rVert_{2}^{2}$ (88)
$\leq-2\eta_{0}\left(-\frac{(\log n)^{c}}{\sqrt{n}}KR_{0}+\frac{1}{2}\lambda_{\min}(\mathbf{\Theta})\right)\lVert\nabla_{\theta}\mathcal{L}^{0}(\theta_{t})\rVert_{2}^{2}$ (89)
@@
Defining $c_{0}:=\frac{1}{3}\lambda_{\min}(\mathbf{\Theta})$ makes it possible to continue with the proof above for $\beta\geq 0$. This is an alternative proof to Lee et al. (2019). It shows that in the unregularized case, it is important that the gradient flow lies in the row space of the Jacobian.
@@ F.2 Closeness to the Linearized Network along the Regularized Gradient Flow
The proof of Lee et al. (2019) used that training error converges to 0. Thus, we have to use a different approach, which also provides a more straightforward and intuitive proof for $\beta=0$. Remember that
@@
$\leq\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}C\eta_{0}KR_{0}e^{-\eta_{0}c_{\beta}t},$ (95)
@@
$\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}(x,\theta_{t})\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}C\eta_{0}KR_{0}\int_{0}^{t}e^{-\eta_{0}c_{\beta}u}du\leq\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}C\frac{KR_{0}}{c_{\beta}}=\frac{(\log n)^{c}}{\sqrt{n}}K^{\prime}C^{2}.$ (96)
In particular, for the difference at the training points, $\lVert f(\theta_{t})-f^{\operatorname*{lin}}(\theta_{t})\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}KC^{2}$.
@@
$\lVert\left(J(\theta_{t})-J(\theta_{0})\right)^{\top}g(\theta_{t})\rVert_{2}\leq\lVert J(\theta_{t})-J(\theta_{0})\rVert_{2}\lVert g(\theta_{t})\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}KCR_{0}.$ (100)
@@
$\lVert J(\theta_{0})^{\top}\left(g(\theta_{t})-g^{\operatorname*{lin}}(\theta_{t})\right)\rVert_{2}=\lVert J(\theta_{0})^{\top}\left(f(\theta_{t})-f^{\operatorname*{lin}}(\theta_{t})\right)\rVert_{2}\leq\lVert J(\theta_{0})\rVert_{2}\lVert f(\theta_{t})-f^{\operatorname*{lin}}(\theta_{t})\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}K^{2}C^{2}.$ (101)
Thus, defining $K^{\Delta}:=KCR_{0}+K^{2}C^{2}$, we can bound $\lVert\Delta_{t}\rVert_{2}\leq\frac{(\log n)^{c}}{\sqrt{n}}K^{\Delta}$. Now, we can compute
$\frac{d}{dt}(\theta_{t}-\theta^{\operatorname*{lin}}_{t})$ $=-\eta_{0}\Delta_{t}-\eta_{0}\left(J(\theta_{0})^{\top}g^{\operatorname*{lin}}(\theta_{t})+\beta(\theta_{t}-\theta_{0})\right)+\eta_{0}\left(J(\theta_{0})^{\top}g^{\operatorname*{lin}}(\theta^{\operatorname*{lin}}_{t})+\beta(\theta^{\operatorname*{lin}}_{t}-\theta_{0})\right)$ (102)
@@
$\leq\int_{0}^{t}e^{-\eta_{0}(c_{0}+\beta)(t-u)}\eta_{0}\frac{(\log n)^{c}}{\sqrt{n}}K^{\Delta}du$ (108)
$\leq\frac{(\log n)^{c}}{\sqrt{n}}\frac{K^{\Delta}}{c_{0}+\beta}.$ (109)
Thus, $\sup_{t}\lVert\theta_{t}-\theta^{\operatorname*{lin}}_{t}\rVert_{2}\leq\frac{K^{\Delta}}{c_{0}+\beta}\frac{(\log n)^{c}}{\sqrt{n}}$.
@@
$\lVert f^{\operatorname*{lin}}(x,\theta_{t})-f^{\operatorname*{lin}}(x,\theta^{\operatorname*{lin}}_{t})\rVert_{2}=\lVert J(x,\theta_{0})(\theta_{t}-\theta^{\operatorname*{lin}}_{t})\rVert_{2}\leq\lVert J(x,\theta_{0})\rVert_{2}\lVert\theta_{t}-\theta^{\operatorname*{lin}}_{t}\rVert_{2}\leq K^{\prime}\frac{K^{\Delta}}{c_{0}+\beta}\frac{(\log n)^{c}}{\sqrt{n}}.$ (110)
By using equation (92), we can now finish the proof:
$\lVert f(x,\theta_{t})-f^{\operatorname*{lin}}(x,\theta^{\operatorname*{lin}}_{t})\rVert_{2}\leq\left(K^{\prime}C^{2}+K^{\prime}\frac{K^{\Delta}}{c_{0}+\beta}\right)\frac{(\log n)^{c}}{\sqrt{n}}.$ (111)